find three consecutive even integers such that four times the second is eight less than the third.
Answer by scott8148(5891)

let x=first, so x+2=second and x+4=third

"four times the second is eight less than the third" __ 4(x+2)+8=x+4

4x+8+8=x+4 __ 3x=-12